OpenStudy (mind2000):

Medals to best answer 52 is 30% of what number

OpenStudy (anonymous):

173.3

OpenStudy (anonymous):

\[\frac{ 30 }{ 100 } / \frac{ 52 }{ x } = \frac{ 5200 }{ 3 } = \frac{ 30x }{ 3 } , x = 173.3\]

OpenStudy (anonymous):

oops typo , divided by 30 , not 3 , still same answer